OpenStudy (noahbred):

Simplify. 15 – (–3) – 4 A. –16 B. –8 C. 22 D. 14

OpenStudy (noahbred):

help

OpenStudy (md152727):

what do you think??

OpenStudy (noahbred):

15

OpenStudy (md152727):

??

OpenStudy (noahbred):

14

OpenStudy (noahbred):

lol

OpenStudy (md152727):

15 is not a multiple choice, yong man

OpenStudy (sloppycanada):

Bing. 14.

OpenStudy (noahbred):

because 15 + 3=18-4 =14

OpenStudy (md152727):

oh :), yep that is right when subtracting in negative numbers we are adding so 15 - -3 means 15+3 or 18 then just 18-4 which is Yes, 14!

OpenStudy (md152727):

Right!

OpenStudy (noahbred):

ty